Two conductors are made of the same material and have the same length. Conductor A is a solid wire of diameter 1.5 mm. Conductor B is a hollow tube of outside diameter 7.4 mm and inside diameter 6.4 mm. What is the resistance ratio RA/RB , measured between their ends?
Answer #1
Concepts and reason

The concepts used to solve this problem are resistance of a solid wire and hollow tube.

First find the resistance for the solid and then find the area of the hollow tube. Later find the resistance for the hollow tube and finally take the ratio of the resistance of these conductors.

Fundamentals

Resistance is defined as the amount of electric current a component can resist.

The expression for resistance is,

R=ρLAR = \frac{{\rho L}}{A}

Here, ρ\rho is the resistivity, LLis the length, and AA is the area.

Express the area of the wire.

A=πd24A = \frac{{\pi {d^2}}}{4}

Here, dd is the diameter of the wire.

Rewrite the expression in terms of area.

R=4ρLπd2R = \frac{{4\rho L}}{{\pi {d^2}}}

The two conductors are made up of the same material, thus, the resistivity is same for both the materials.

The expression for resistance ratio is,

Resistanceratio=RARB{\rm{Resistance}}\,{\rm{ratio}} = \frac{{{R_A}}}{{{R_B}}}

Here, RA{R_A} is the resistance of a solid wire and RB{R_B} is the resistance of a hollow tube.

The expression for resistance of the solid wire is,

RA=4ρLπd2{R_A} = \frac{{4\rho L}}{{\pi {d^2}}}

Substitute 1.5mm1.5\,{\rm{mm}} for the diameter dd.

RA=4ρLπ(1.5mm×103m1mm)2=ρL1.76×106m2\begin{array}{c}\\{R_A} = \frac{{4\rho L}}{{\pi {{\left( {{\rm{1}}{\rm{.5}}\,{\rm{mm }} \times \frac{{{{10}^{ - 3}}\,{\rm{m}}}}{{1\,{\rm{mm}}}}} \right)}^2}}}\\\\ = \frac{{\rho L}}{{1.76 \times {{10}^{ - 6}}{\rm{ }}{{\rm{m}}^2}}}\\\end{array}

The expression to calculate the resistance of the hollow tube is,

RB=4ρLπd2{R_B} = \frac{{4\rho L}}{{\pi {d^2}}}

The expression to calculate the area of the hollow tube is,

πd24=π4(douter2dinner2)\frac{{\pi {d^2}}}{4} = \frac{\pi }{4}\left( {d_{{\rm{outer}}}^2 - d_{{\rm{inner}}}^2} \right)

Here, douter{d_{{\rm{outer}}}} is the outer diameter of the hollow tube and dinner{d_{{\rm{inner}}}} is the inner diameter of the hollow tube.

Use π4(douter2dinner2)\frac{\pi }{4}\left( {d_{{\rm{outer}}}^2 - d_{{\rm{inner}}}^2} \right) for πd24\frac{{\pi {d^2}}}{4} in RB=4ρLπd2{R_B} = \frac{{4\rho L}}{{\pi {d^2}}}. The resistance of the hollow tube is,

RB=4ρLπ(douter2dinner2){R_B} = \frac{{4\rho L}}{{\pi \left( {d_{{\rm{outer}}}^2 - d_{{\rm{inner}}}^2} \right)}}

Substitute 7.4 mm for douter{d_{{\rm{outer}}}} and 6.4 mm for dinner{d_{{\rm{inner}}}} in the above expression.

RB=4ρLπ(((7.4mm)(1m103mm))2((6.4mm)(1m103mm))2)=ρL1.0833×105m2\begin{array}{c}\\{R_B} = \frac{{4\rho L}}{{\pi \left( {{{\left( {\left( {7.4\,{\rm{mm}}} \right)\left( {\frac{{1\,{\rm{m}}}}{{{{10}^3}\,{\rm{mm}}}}} \right)} \right)}^2} - {{\left( {\left( {6.4\,{\rm{mm}}} \right)\left( {\frac{{1\,{\rm{m}}}}{{{{10}^3}\,{\rm{mm}}}}} \right)} \right)}^2}} \right)}}\\\\ = \frac{{\rho L}}{{1.0833 \times {{10}^{ - 5}}\,{{\rm{m}}^{\rm{2}}}}}\\\end{array}

Resistanceratio=RARB{\rm{Resistance}}\,{\rm{ratio}} = \frac{{{R_A}}}{{{R_B}}}

Substitute ρL1.76×106m2\frac{{\rho L}}{{1.76 \times {{10}^{ - 6}}{\rm{ }}{{\rm{m}}^2}}} for RA{R_A} and [ρL1.08×105m2]\left[ {\frac{{\rho L}}{{1.08 \times {{10}^{ - 5}}{\rm{ }}{{\rm{m}}^2}}}} \right] for RB{R_B} inRARB\frac{{{R_A}}}{{{R_B}}}.

RARB=(ρL1.76×106m2)(ρL1.08×105m2)=[ρL1.76×106m2][1.08×105m2ρL]=1.08×105m21.76×106m2=6.136\begin{array}{c}\\\frac{{{R_A}}}{{{R_B}}} = \frac{{\left( {\frac{{\rho L}}{{1.76 \times {{10}^{ - 6}}{\rm{ }}{{\rm{m}}^2}}}} \right)}}{{\left( {\frac{{\rho L}}{{1.08 \times {{10}^{ - 5}}{\rm{ }}{{\rm{m}}^2}}}} \right)}}\\\\ = \left[ {\frac{{\rho L}}{{1.76 \times {{10}^{ - 6}}{\rm{ }}{{\rm{m}}^2}}}} \right]\left[ {\frac{{1.08 \times {{10}^{ - 5}}{\rm{ }}{{\rm{m}}^2}}}{{\rho L}}} \right]\\\\ = \frac{{1.08 \times {{10}^{ - 5}}{\rm{ }}{{\rm{m}}^2}}}{{1.76 \times {{10}^{ - 6}}{\rm{ }}{{\rm{m}}^2}}}\\\\ = 6.136\\\end{array}

Ans:

The resistance ratio RARB\frac{{{R_A}}}{{{R_B}}} measured between their ends is 6.136.